### 19 Apr 2026 v2

# Lightweight LLM Agent Memory with Small Language Models

Abstract
Although LLM agents can leverage tools for complex tasks, they still need memory to maintain cross-turn consistency and accumulate reusable information in long-horizon interactions. However, retrieval-based external memory systems incur low online overhead but suffer from unstable accuracy due to limited query construction and candidate filtering. In contrast, many systems use repeated large-model calls for online memory operations, improving accuracy but accumulating latency over long interactions. We propose LightMem, a lightweight memory system for better agent memory driven by Small Language Models (SLMs). LightMem modularizes memory retrieval, writing, and long-term consolidation, and separates online processing from offline consolidation to enable efficient memory invocation under bounded compute. We organize memory into short-term memory (STM) for immediate conversational context, mid-term memory (MTM) for reusable interaction summaries, and long-term memory (LTM) for consolidated knowledge, and uses user identifiers to support independent retrieval and incremental maintenance in multi-user settings. Online, LightMem operates under a fixed retrieval budget and selects memories via a two-stage procedure: vector-based coarse retrieval followed by semantic consistency re-ranking. Offline, it abstracts reusable interaction evidence and incrementally integrates it into LTM. Experiments show gains across model scales, with an average F1 improvement of about 2.5 on LoCoMo, more effective and low median latency (83 ms retrieval; 581 ms end-to-end).
Lightweight LLM Agent Memory with Small Language Models
Jiaquan Zhang1, Chaoning Zhang1,*, Shuxu Chen2, Zhenzhen Huang1, Pengcheng Zheng1, Zhicheng Wang1, Ping Guo3, Fan Mo4, Sung-Ho Bae2, Jie Zou1, Jiwei Wei1, Yang Yang1 1University of Electronic Science and Technology of China; 2Kyung Hee University; 3City University of Hong Kong; 4University of Oxford *Corresponding author: chaoningzhang1990@gmail.com
1 Introduction
LLM-driven agents excel at long-term dialogue, multi-step reasoning, and task-oriented interaction Zhang et al. (2026c); Huang et al. (2024); Zhu et al. (2025); Zhang et al. (2026a, b, 2025a). To maintain cross-turn consistency beyond the context window, many systems augment agents with external memory Lee et al. (2024); Xu et al. (2025); Hu et al. (2025); Wang et al. (2026). Long-term memory supports continual learning and planning.
Figure 1: LightMem combines enhanced retrieval with SLMs, achieving high retrieval accuracy while significantly reducing online latency compared to retrieval-based and LLM-based memory systems.
Existing memory systems can be broadly divided into two categories. One line is based on retrieval–driven external memory, where past interactions are compressed into retrievable entries Asai et al. ; Maharana et al. (2024) and the top-K memories are recalled via similarity search Zhong et al. (2024). These methods are efficient, but limited query construction and candidate filtering often introduce retrieval noise, resulting in unstable answer accuracy. Another line of methods Tan et al. (2025); Zhang et al. (2025b) introduces LLM-driven memory operations on top of external memory, repeatedly invoking large models for memory writing, retrieval and controlling. As these operations are typically implemented via repeated model invocations, they can introduce non-trivial runtime overhead over long interactions Packer et al. (2023); Park et al. (2023). This contrast suggests a natural separation: keep high-frequency online memory decisions lightweight and controllable, while deferring heavy abstraction and consolidation to offline processing (see Figure 1). Recent advances in SLMs Magister et al. (2023); Han et al. (2024) make this separation practical. SLMs can reliably handle high-frequency, structured decision tasks in online memory processing, enabling different memory stages to be assigned to models of appropriate scale Liu et al. (2025). Online memory handling typically consists of structured subtasks (such as intent routing Zhang et al. (2024), query construction Hong and He (2025) and semantic filtering Hatalis et al. (2023)). These tasks place greater emphasis on predictable behavior and low overhead than on maximal generative capacity. In this setting, lightweight SLMs provide a suitable choice Sinha et al. (2025) for online control and filtering, while heavier abstraction and consolidation can be deferred to offline processing. Overall, existing memory systems face an efficiency–effectiveness trade-off. Moreover, SLMs are not a silver bullet due to limited capacity and representation. Bridging this gap requires strengthening the external memory pipeline, especially how memories are written and retrieved.
To address the above questions, we propose LightMem, a lightweight memory system for LLM agents that models long-term memory as an incrementally evolvable process via specialized SLMs. LightMem modularizes and decouples query parsing, memory retrieval, memory writing, and long-term consolidation, enabling independent optimization and elastic scaling of each component. This separation allows for lightweight online processing to be distinct from offline consolidation, achieving efficient memory use under tight compute budgets and supporting long-term evolution. Memories are organized into STM, MTM, and LTM stores based on temporal and access characteristics. User-identity metadata is embedded in each memory unit to enforce user-level logical isolation, balancing privacy, consistency, and scalability. We propose a modular online–offline memory pipeline in which distinct SLMs specialize in complementary memory operations. Online, LightMem uses three specialized small language model modules: a Controller (SLM-1) for intent and query planning, a Selector (SLM-2) for candidate verification and compression, and a Writer (SLM-3) for incremental memory writing. Specifically, the Controller rewrites the user input into intent-conditioned hypothetical queries (HQs) and allocates a fixed Top-$K$ budget across MTM and LTM. Given the retrieved candidates, the Selector performs metadata-constrained prefiltering followed by semantic-consistency based re-ranking to output the final Top-$K$ memories. After each turn, the Writer summarizes the interaction into compact MTM entries and maintains MTM incrementally. Long-term abstraction and consolidation are handled offline by a large-context model, keeping the online path lightweight. Our main contributions are summarized as:
  • We propose LightMem, a lightweight memory system collaboratively driven by SLMs. Different SLMs handle lightweight, high-frequency online memory operations (query construction, retrieval, and writing), while heavier abstraction and consolidation are deferred to offline processing.
  • We propose a two-stage memory querying design. Given a user query, LightMem first narrows down the candidate set with fast retrieval, and then applies semantic-level verification to select the truly relevant memories.
  • We evaluate LightMem on LoCoMo and DialSim, demonstrating gains across model scales, including an average F1 improvement of about 2.5 on LoCoMo, improved semantic consistency on DialSim and low median latency (83 ms retrieval; 581 ms end-to-end).
2 Related Work
2.1 LLM Memory Systems
Prior work can be broadly categorized into retrieval-based memory and LLM-driven memory operations. Retrieval-based Memory. MemoryBank Zhong et al. (2024) supports personalized long-term dialogue by storing summarized user events in an external memory and retrieving relevant entries for each query, with forgetting to control growth. MemGPT Packer et al. (2023) treats the context window as virtual memory and performs paging between the prompt and an external store, with runtime eviction and on-demand retrieval. ReadAgent Lee et al. (2024) similarly relies on retrieval over an external cache: it indexes compressed gists and performs on-demand lookup to fetch supporting evidence when needed. LLM-driven Memory. HiAgent Hu et al. (2025) manages in-trial working memory by chunking trajectories into subgoals and summarizing past steps with LLMs. A-MEM Xu et al. (2025) further builds self-organizing memory networks through LLM-driven note-taking and automatic linking, but typically does not emphasize strict online/offline decoupling under resource constraints. These lines expose a recurring trade-off between lightweight but noisy retrieval-based memory and more effective yet costly LLM-driven online memory operations. LightMem addresses it with SLM-based lightweight online control and filtering, coupled with improved external memory writing and retrieval under a fixed budget.
3 Method
As shown in Figure 2, LightMem uses specialized SLMs to modularize memory operations, separating lightweight online querying from offline long-term consolidation.
Figure 2: Multiple SLMs coordinate an online pathway for query-time routing and retrieval over STM/MTM, and an offline pathway that incrementally consolidates MTM into a graph-structured LTM.
3.1 Problem Setup and Preliminaries
We consider a multi-turn dialogue setting with multiple users. At turn $t$, a user provides an input $x_{t}$, and the model generates a response $y_{t}$. Due to the limited context capacity of SLMs, the accessible dialogue context at turn $t$ is denoted by $C_{t}$, typically a truncated window of recent turns. LightMem maintains a user-scoped memory store $M_{u}$ with user-level logical isolation via user identifiers (see §3.2 for the definition of STM/MTM/LTM). Given $(x_{t},C_{t})$, the system retrieves a memory set $R_{t}\subseteq M_{u}$ with $|R_{t}|\leq K$ and uses $R_{t}$ to assist in generating $y_{t}$. The specific symbols are provided in the Appendix 8.1.
3.2 Memory Stores
STM.
STM is implemented as the SLM context window that buffers the most recent interaction sequence in the current session. It serves only as working memory and is updated turn by turn within the prompt; STM is neither persisted nor retrieved.
MTM.
MTM is the sole carrier of personalized episodic memory. It stores a set of memory items, each consisting of (i) a concise semantic summary, (ii) temporal information and access statistics, (iii) an embedding for similarity-based retrieval, and (iv) a user identifier for strict per-user isolation.
LTM.
LTM stores de-identified, user-agnostic semantic knowledge distilled offline from high-value MTM episodes. It contains no raw personal episodes or user identifiers; instead, it represents stable facts, domain regularities, and cross-user trends. LTM is organized as a lightweight graph-structured knowledge base to support multi-hop reasoning and knowledge sharing across tasks.
3.3 Intent Modeling and Retrieval Control
Given $(x_{t},C_{t})$, SLM-1 serves as a lightweight retrieval controller rather than a retriever. It converts the raw input into a structured retrieval plan that specifies: (i) what to search (queries), (ii) how to constrain the search (metadata filters), and (iii) how many items to return under a fixed Top-$K$ budget. Concretely, SLM-1 first infers coarse intent attributes, e.g., whether the query depends more on recent episodic details or long-term stable knowledge, and whether strong personalization is needed. These attributes are used only for retrieval planning (query rewriting, routing, and budgeting), not for answer generation. SLM-1 then rewrites the input into a set of HQs $\{q_{t}^{(i)}\}$ and outputs metadata constraints $\phi_{t}$ (user identifier, optional time window, and type tags) to reduce noise and enforce user-level isolation. Finally, it issues the retrieval request:
$\mathcal{Q}_{t}=\left\langle\{q_{t}^{(i)}\},\ \phi_{t},\ K\right\rangle.$ (1)
This request is consumed by §3.4.
3.4 Two-Stage Retrieval
Given $\mathcal{Q}_{t}$, SLM-2 executes a two-stage retrieval procedure and returns a memory set $R_{t}$ that satisfies the Top-$K$ constraint:
$R_{t}=\textsc{Retrieve}(\mathcal{Q}_{t}),\qquad|R_{t}|\leq K.$ (2)
The retrieved memories $R_{t}$ are appended to the context to condition response generation for $y_{t}$.
For a single user query, SLM-1 may generate $n$ HQs, denoted by $\{q_{t}^{(i)}\}_{i=1}^{n}$, to improve recall coverage. We impose a single final return budget $K$ and adopt a double-budget rule in the coarse stage: the total number of candidates returned by Stage 1 is fixed to $2K$.
Stage 1: metadata-constrained coarse retrieval.
Under metadata constraints $\phi_{t}$ (including the user identifier, an optional time window, and type/source tags), the system performs vector-based coarse retrieval for each HQ. The Stage 1 candidate budget is evenly split across HQs:
$\sum_{i=1}^{n}K_{1}^{(i)}=2K,\qquad K_{1}^{(i)}=\frac{2K}{n}.$ (3)
Let $C^{(i)}$ denote the candidate list returned for the $i$-th HQ. The aggregated candidate set is
$C=\bigcup_{i=1}^{n}C^{(i)},\qquad|C|=2K.$ (4)
Stage 1 is designed for coverage, acting as an efficient filter that compresses the large search space into a fixed-size candidate set.
Stage 2: semantic filtering and compression.
In Stage 2, SLM-2 takes the HQ set $\{q_{t}^{(i)}\}_{i=1}^{n}$ and the Stage 1 candidates $C$ (memory summaries with necessary structured metadata) and performs semantic consistency checking and relevance judgment. Instead of generating answers, SLM-2 conducts controlled selection on a fixed-size candidate pool: it keeps the most relevant half of $|C|=2K$ candidates and outputs the final retrieval results
$R_{t}\subseteq C,\qquad|R_{t}|\leq K$ (5)
This two-to-one compression yields (i) stable computation with a fixed candidate size, (ii) semantic refinement beyond vector similarity by leveraging intent and metadata, and (iii) noise suppression by explicitly discarding roughly half of the candidates. The resulting Top-$K$ memories $R_{t}$ are provided to the downstream agent or response generator.
3.5 Memory Writing and Update
After generating $y_{t}$, SLM-3 extracts reusable user-relevant information from the current interaction, compresses it into a concise memory item, and appends it to MTM. To prevent redundancy and noise accumulation, highly repetitive items are merged or rewritten, while conflicting information is handled using temporal cues and evidence strength. To ensure consistently low-latency personalized retrieval, we enforce a capacity bound on MTM:
$|M_{u}^{\textsc{MTM}}|\leq B$ (6)
When the bound is reached, MTM is maintained by evicting stale, low-utility items and further compressing redundant content.
3.6 Offline Consolidation
To avoid increasing online retrieval and writing latency, LightMem performs offline consolidation to distill high-value episodic evidence in MTM into de-identified, long-term semantic knowledge, enabling sustained evolution of LTM. An LLM handles this offline path with a large context window and is strictly decoupled from online operation. In each cycle, the LLM processes only an incremental batch (newly written or retrieval-reactivated MTM items and low-utility candidates flagged under MTM capacity pressure), rather than rebuilding MTM/LTM from scratch. It abstracts episodes into privacy-preserving knowledge candidates, performs similarity search over LTM to locate semantically nearest anchors, and incrementally inserts and links candidates within the local neighborhood, thereby maintaining LTM as lightweight graph-structured knowledge. Evidence accumulated over time further drives merge/update/drop decisions, while confidence decay is applied to weakly supported candidates to enable natural forgetting and to limit the impact of stale or incidental information on subsequent retrieval and reasoning. We provide each component with algorithmic details deferred to Appendix 8.1.
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
To evaluate the effectiveness of LightMem, we conduct experiments on two datasets. LoCoMo Maharana et al. (2024) is a benchmark designed to evaluate logical reasoning over extended conversational contexts. It contains long dialogues (on average 9K tokens) and covers five core task categories: Single-hop, Multi-hop, Temporal, Open-domain, and Adversarial. DialSim Kim et al. (2024) is a dialogue simulation dataset derived from TV shows, containing 1,300 multi-party sessions spanning years.
4.1.2 Baselines
Baselines We select baselines that are representative in prior work on long-term memory, covering common designs for storing, retrieving, and updating dialogue histories and that can be evaluated under the same backbone and retrieval budget for a fair comparison: LoCoMo Maharana et al. (2024), ReadAgent Lee et al. (2024), MemoryBank Zhong et al. (2024), MemGPT Packer et al. (2023), and A-MEM Xu et al. (2025). When selecting backbone models, a diverse array of LLMs is used to assess their generalization capabilities. These models encompass GPT-4o, GPT-4o-mini Hurst et al. (2024), Qwen2.5 (available in 1.5B and 3B) Bai et al. (2025), and Llama 3.2 (offered in 1B and 3B) Dubey et al. (2024).
4.1.3 Metric
We evaluate response quality using both lexical-overlap and semantic metrics. Specifically, we report F1 and BLEU-1 to measure answer correctness and token-level overlap with reference responses, which are commonly used for short factual and span-like outputs. For DialSim Kim et al. (2024), where multiple surface forms can express the same meaning, we additionally report ROUGE-L and METEOR to capture sequence-level overlap and soft token matching, and use SBERT similarity to quantify semantic consistency between generated and reference responses beyond n-gram matches. These metrics provide a reliable assessment of both precision and meaning retention in long-horizon dialogue scenarios.
4.1.4 Implement Details
To isolate the effect of memory, we use GPT-4o-mini as the fixed response generator for LightMem and all baselines. LightMem’s control plane uses locally deployed, quantized SLMs: SLM-1 for HQ generation and retrieval control, SLM-2 for semantic re-ranking and compression, and SLM-3 for online writing and MTM maintenance. For HQ generation, SLM-1 follows a structured prompt for query decomposition and routing; Table 1 summarizes the prompt logic and provides an example. We deploy quantized Llama-3.2-1B-Instruct via Ollama by default and additionally evaluate Qwen2.5-1.5B-Instruct for robustness. Offline consolidation is handled by a large-context LLM and is decoupled from the online path. We encode all queries and memory entries using all-MiniLM-L6-v2 (384 dimensions) and retrieve the top-10 nearest neighbors in the coarse vector retrieval stage. SLM-2 is fine-tuned with LoRA on 2,000 constructed (Query, Subgraph, Path) samples. We cap MTM capacity at $B=10^{4}$ and apply dynamic pruning beyond this limit by evicting stale, low-utility items and merging/compressing redundant content. LightMem inference and latency measurements are conducted on a single NVIDIA RTX 4090 (24GB) GPU. Other unspecified settings follow A-MEM Xu et al. (2025).
Table 1: Structured HQ prompt used by SLM-1 for query decomposition and routing.
HQ prompt step Example
Detect missing info Identify underspecified references (e.g., pronouns such as “it/that/he/the project”), implicit context dependencies, and vague time cues (e.g., “recently/last time/before”).
Generate HQs Rewrite the user request into one or more standalone HQs, optionally splitting intent into user-specific vs. factual queries, and making implied context (time/location) explicit.
Route and budget Assign each HQ to MTM (user-specific) or LTM (general/public knowledge) and allocate a retrieval budget under a fixed top-$K$ constraint.
Example Input: “Recommend a dinner spot.” The prompt yields two HQs: one asks about the user’s preferred cuisines or dietary constraints (routed to MTM), and the other asks about highly rated nearby restaurants (routed to LTM).
Table 2: Main results on LoCoMo across question categories. We report F1 and BLEU-1 for each category, along with the token length of the effective context. Best results within each model block are in bold.
Model Method Single-hop Multi-hop Temporal Open-domain Adversarial Token Length
F1 BLEU F1 BLEU F1 BLEU F1 BLEU F1 BLEU
GPT-4o-mini LoCoMo 40.36 29.05 25.02 19.75 18.41 14.77 12.04 11.16 69.23 68.75 16,910
ReadAgent 9.67 7.66 9.15 6.48 12.60 8.87 5.31 5.12 9.81 9.02 643
MemoryBank 6.61 5.16 5.00 4.77 9.68 6.99 5.56 5.94 7.36 6.48 432
MemGPT 41.04 34.34 26.65 17.72 25.52 19.44 9.15 7.44 43.29 42.73 16,977
A-MEM 44.65 37.06 27.02 20.09 45.85 36.67 12.14 12.00 50.03 49.47 2,520
LightMem 45.80 38.20 28.85 21.40 46.20 37.10 13.50 12.90 54.50 52.10 1,150
GPT-4o LoCoMo 61.56 54.19 28.00 18.47 9.09 5.78 16.47 14.80 52.61 51.13 16,910
ReadAgent 12.46 10.29 14.61 9.95 4.16 3.19 8.84 8.37 6.81 6.13 805
MemoryBank 8.28 7.10 6.49 4.69 2.47 2.43 6.43 5.30 4.42 3.67 569
MemGPT 60.16 53.35 30.36 22.83 17.29 13.18 12.24 11.87 34.96 34.25 16,987
A-MEM 48.43 42.97 32.86 23.76 39.41 31.23 17.10 15.84 36.35 35.53 1,216
LightMem 49.80 44.20 34.50 25.10 40.10 32.00 18.40 16.90 41.50 39.80 680
Qwen2.5-1.5B LoCoMo 11.15 8.67 9.05 6.55 4.25 4.04 9.91 8.50 40.38 40.23 16,910
ReadAgent 10.13 7.54 6.61 4.93 2.55 2.51 5.31 12.24 5.42 27.32 752
MemoryBank 13.42 11.01 11.14 8.25 4.46 2.87 8.05 6.21 36.76 34.00 284
MemGPT 9.56 7.34 10.44 7.61 4.21 3.89 13.42 11.64 31.51 28.90 16,953
A-MEM 23.63 19.23 18.23 11.94 24.32 19.74 16.48 14.31 46.00 43.26 1,300
LightMem 25.10 20.50 21.50 13.80 25.60 20.80 17.20 15.10 49.80 46.50 720
Qwen2.5-3B LoCoMo 7.03 5.69 4.61 4.29 3.11 2.71 4.55 5.97 16.95 14.81 16,910
ReadAgent 3.25 2.51 2.47 1.78 3.01 3.01 5.57 5.22 15.78 14.01 776
MemoryBank 4.11 3.32 3.60 3.39 1.72 1.97 6.63 6.58 13.07 10.30 298
MemGPT 7.26 5.52 5.07 4.31 2.94 2.95 7.04 7.10 14.47 12.39 16,961
A-MEM 17.23 13.12 12.57 9.01 27.59 25.07 7.12 7.28 27.91 25.15 1,137
LightMem 19.80 15.20 15.80 11.20 29.10 26.50 9.50 9.10 33.50 29.80 610
Llama-3.2-1B LoCoMo 12.86 10.50 11.25 9.18 7.38 6.82 11.90 10.38 51.89 48.27 16,910
ReadAgent 7.75 6.03 5.96 5.12 1.93 2.30 12.46 11.17 44.64 40.15 665
MemoryBank 17.30 14.03 13.18 10.03 7.61 6.27 15.78 12.94 52.61 47.53 274
MemGPT 10.16 7.68 9.19 6.96 4.02 4.79 11.14 8.24 49.75 45.11 16,950
A-MEM 28.51 24.13 19.06 11.71 17.80 10.28 17.55 14.67 58.81 54.28 1,376
LightMem 30.80 26.20 22.40 14.20 19.90 12.10 18.90 16.10 63.50 58.40 750
Llama-3.2-3B LoCoMo 8.37 6.93 6.88 5.77 4.37 4.40 10.65 9.29 30.25 28.46 16,910
ReadAgent 3.25 2.51 2.47 1.78 3.01 3.01 5.57 5.22 15.78 14.01 461
MemoryBank 7.61 6.03 6.19 4.47 3.49 3.13 4.07 4.57 18.65 17.05 263
MemGPT 4.32 3.51 5.32 3.99 2.68 2.72 5.64 5.54 21.45 19.37 16,956
A-MEM 28.14 23.87 17.44 11.74 26.38 19.50 12.53 11.83 42.04 40.60 1,126
LightMem 30.20 25.90 20.10 13.50 28.20 21.10 14.20 13.50 47.10 44.80 640
4.2 Performance
Performance. As shown in Table 2, LightMem achieves the best overall performance on the LoCoMo benchmark. In comparison to baselines that directly rely on long-context replay (such as LoCoMo and MemGPT), LightMem demonstrates more stable performance across the majority of question categories. It has particularly distinct advantages for multi-hop and temporal questions, all while not relying on a single long-context input. In comparison with summarization- or compression-based methods (MemoryBank and ReadAgent), LightMem consistently outperforms them across all categories, with larger margins on multi-hop and temporal reasoning. When compared to the strongest memory-centric baseline, A-MEM, LightMem further improves overall performance in most model settings. On GPT-4o, LightMem attains an F1 score of 34.50 for multi-hop questions, surpassing A-MEM’s score of 32.86. Overall, LightMem demonstrates more robust long-term memory across different model scales and question types. In addition, it maintains significant advantages in cross-session reasoning tasks.
Table 3 reports the comparison on DialSim with GPT-4o-mini. We incorporate METEOR and SBERT similarity. This is driven by the characteristics of long-term dialogue generation. In long-term dialogue generation, semantically accurate answers can be presented in diverse surface forms. Consequently, metrics based solely on n-grams may undervalue advancements. METEOR partially accounts for morphological and synonym-level matching, while SBERT directly measures semantic alignment between the generated response and the reference, providing a complementary view of memory effectiveness. Across all metrics, LightMem achieves the best results. Importantly, LightMem improves not only lexical overlap (e.g., ROUGE-L/ROUGE-2 and METEOR) but also semantic similarity, with SBERT increasing from 19.51 (A-MEM) to 23.40. This indicates that the gains are not limited to reproducing similar wording, but also reflect stronger semantic consistency enabled by more effective use of long-term conversational history.
Table 3: Comparison of different memory mechanisms on DialSim with GPT-4o-mini.
Method F1 BLEU-1 ROUGE-L ROUGE-2 METEOR SBERT
LoCoMo 2.55 3.13 2.75 0.90 1.64 15.76
MemGPT 1.18 1.07 0.96 0.42 0.95 8.54
A-MEM 3.45 3.37 3.54 3.60 2.05 19.51
LightMem 4.12 3.95 4.20 4.15 2.48 23.40
Generalization. We evaluate all methods across diverse backbones, ranging from large models (GPT-4o / GPT-4o-mini) to small open-source models (Qwen2.5 and Llama-3.2). As shown in Table 2, LightMem remains the best or near-best method across these settings, indicating that its gains are robust and not tied to a specific backbone.
4.3 Ablation Study
To quantify the contribution of each component in LightMem, we conduct ablation studies by removing or simplifying one key module at a time while keeping the rest of the system unchanged. We consider the following variants:
  • w/o semantic reranking We remove the LM-based semantic filtering stage and keep only embedding-based Top-$K$ retrieval.
  • w/o HQ and retrieval routing We disable HQs as well as the controller-based routing/budget allocation, and retrieve using the original query directly.
  • w/o MTM We remove the MTM layer, relying only on the current context (STM) and LTM for retrieval and generation.
  • w/o offline consolidation We disable the offline consolidation pipeline so that LTM is not periodically updated from MTM.
  • w/o graph structure We replace the graph-structured LTM with a flat vector store.
(a) Ablation study on Dialsim with Llama-3.2-1B
(b) Ablation study on Dialsim with Qwen2.5-1.5B
Figure 3: Ablation study on DialSim. We report F1, BLEU-1, ROUGE-L, ROUGE-2, METEOR, and SBERT similarity.
Figure 3 shows that removing any single component consistently degrades performance across metrics. Disabling semantic reranking or HQ-based retrieval routing leads to clear drops in F1; on Llama-3.2-1B, LightMem achieves an F1 of 4.12, while the variant without semantic reranking scores 3.83 and the variant without HQ and retrieval routing scores 3.87. Removing MTM further lowers performance, with an F1 of 3.75, indicating the benefit of retaining mid-term episodic memory. In comparison, disabling offline consolidation results in a smaller but still consistent decline, with an F1 of 3.96 on Llama-3.2-1B. Removing the graph structure also reduces both lexical and semantic quality, as reflected by a lower SBERT similarity of 22.82 compared to 23.40 for the full LightMem. Overall, the full LightMem achieves the best results, suggesting that these components contribute complementary benefits.
Table 4: Scalability and latency analysis across five baselines on GPT-4o-mini.
[…]
LoCoMo 0 0 2054 3658
ReadAgent 34 97 614 1027
MemoryBank 16 51 439 1047
MemGPT 143 451 2087 3451
A-MEM 856 1583 914 3682
LightMem 83 167 581 1325
4.4 Scalability and Latency
Under a unified deployment configuration, we assess the scalability and latency by utilizing GPT-4o-mini. Subsequently, we present the following latency metrics:
  • Retrieval Latency (ms). Time from receiving a user query to finishing memory retrieval and constructing the final prompt. It includes local computation but excludes remote GPT-4o-mini generation time. For the full-context LoCoMo baseline, retrieval latency is 0 ms since no external memory is retrieved.
  • End-to-End Latency (ms). The time from receiving a user query to producing the final response, including retrieval, remote API transmission and queuing, and GPT-4o-mini inference and decoding.
  • P50 / P95. The median (P50) and 95th-percentile (P95) latency across all samples. P95 captures tail latency from network jitter and service queuing, reflecting real-world system stability.
Table 4 compares retrieval and end-to-end latency on GPT-4o-mini. Among memory-enabled methods, LightMem achieves a good balance between retrieval overhead and end-to-end responsiveness. Its retrieval latency remains low, with P50/P95 of 83/167 ms, substantially lower than MemGPT and especially A-MEM, which exhibits heavy tail latency. LightMem also shows competitive end-to-end latency, with a P50 of 581 ms and a controlled P95, indicating stable behavior under remote API settings. When considered together with the effective context length in Table 2, these results are more interpretable: LoCoMo and MemGPT rely on replaying large dialogue histories with effective contexts close to 16K tokens, whereas LightMem consistently operates with much shorter contexts (around 1K tokens on GPT-4o-mini). This reduction in effective context length lowers prompt construction and inference cost, contributing to better scalability and practical deployability.
4.5 Performance Stability under MTM Growth
We analyze how MTM size affects performance stability on DialSim with Llama-3.2-1B. Rather than rerunning the model with different memory sizes, we report cumulative F1 at several natural growth stages along the same full dialogue trajectory, approximately at 100, 1,000, 5,000, and 10,000 MTM entries. As shown in Table 5, the gap between the two methods is small when MTM remains relatively small. However, as memory entries accumulate along the dialogue trajectory, pure vector retrieval becomes increasingly affected by retrieval noise, whereas LightMem maintains more stable performance. This trend is consistent with the role of Stage 2 semantic filtering, which helps preserve retrieval quality as the memory store grows.
Table 5: Performance stability under natural MTM growth on DialSim with Llama-3.2-1B. Statistics are computed cumulatively along the same full dialogue trajectory at different MTM sizes.
[…]
$\sim$100 3.95 3.98 0.03
$\sim$1,000 3.90 4.05 0.15
$\sim$5,000 3.86 4.09 0.23
10,000 3.83 4.12 0.29
4.6 Error Injection Stress Test
To study error accumulation in long-horizon memory systems, we conduct an end-to-end error injection stress test on DialSim under a fixed final Top-$K$ memory budget. Group A is the full LightMem system. Group B injects 50% irrelevant HQs to simulate SLM-1 failures in query planning. Group C removes SLM-2 semantic reranking and directly uses the stage-one vector Top-$K$. Group D replaces 50% of SLM-3 writes with noisy strings, simulating persistent MTM pollution. Group E combines all three perturbations to simulate cascading failure. As shown in Table 6, performance degrades consistently as errors are introduced into different stages of the pipeline. Query noise in SLM-1 causes a moderate drop, suggesting partial robustness to retrieval-side perturbations. Removing SLM-2 leads to a larger decline, confirming the importance of semantic filtering after coarse retrieval. Noisy writes in SLM-3 further degrade performance, indicating that memory pollution introduces persistent downstream effects by contaminating future retrieval. Cascading failure causes the largest collapse, showing that errors can accumulate and amplify across turns when noisy retrieval, missing filtering, and corrupted writing occur simultaneously.
Table 6: Error injection stress test on DialSim under a fixed final Top-$K$ memory budget.
Group Setting F1 SBERT
A Full LightMem 4.12 23.40
B SLM-1 50% HQ noise 3.95 23.08
C w/o SLM-2 reranking 3.83 22.56
D SLM-3 50% write noise 3.78 22.45
E Cascading failure 1.85 11.20
4.7 Update Gap Stress Test
As shown in Table 7, concurrent retrieval from MTM and LTM is important for handling the update gap. The full LightMem system achieves the best performance, with a multi-hop F1 of 28.85, whereas using only LTM or only MTM results in clear performance drops. This suggests that concurrent routing effectively bridges recent unconsolidated facts in MTM and older consolidated knowledge in LTM. A more challenging case appears in Group D. Under MTM noise saturation, the multi-hop F1 decreases to 23.10, indicating that heavy recent noise can interfere with retrieval during the update gap. This is likely because the fixed Top-$K$ budget is increasingly occupied by irrelevant recent records, making the newest relevant fact harder to retrieve before it is consolidated into LTM. Overall, the results show that concurrent retrieval provides a practical and effective solution for update-gap reasoning, especially when recent and long-term evidence must be combined.
Table 7: Update gap stress test on the LoCoMo multi-hop subset with GPT-4o-mini.
Group Setting Multi-hop F1
A Full LightMem (normal gap) 28.85
B LTM-only retrieval 19.45
C MTM-only retrieval 20.12
D MTM noise saturation 23.10
5 Conclusion
We present LightMem, a lightweight external memory system for LLM agents driven by SLMs. LightMem decouples high-frequency online memory operations (query control, retrieval, and writing) from offline consolidation, and organizes memory into STM/MTM/LTM with user-scoped isolation. Online, it operates under a fixed retrieval budget and uses HQ-based routing plus a two-stage retrieval (coarse vector search followed by semantic consistency re-ranking) to reduce retrieval noise. Experiments on LoCoMo and DialSim show consistent gains across model scales, including  +2.5 average F1 on LoCoMo, improved semantic consistency on DialSim, and low median latency (83 ms retrieval; 581 ms end-to-end).
6 Limitations
This work focuses on a specific design of online–offline decoupled memory pipelines driven by specialized small language models. The impact of alternative consolidation strategies and control policies is not fully explored and remains an interesting direction for future investigation.
7 Acknowledgements
This work was partially supported by the National Natural Science Foundation of China under grant 62572104, and 62220106008.